OpenStudy (anonymous):

Given the parabola f(x)=x^2+4x+21, state and Identify the max or min value.

OpenStudy (kira_yamato):

f(x) = x^2 + 4x + 21 f'(x) = 2x + 4 f''(x) = 2 > 0 (min) For min value, f'(x) = 0 2x = -4 x = -2 f(-2) = 4 - 8 + 21 = 17

OpenStudy (anonymous):